Your cable bill is $122 per month (starting next month) If your investments earn 4.00% APR (compounded monthly) how much would you have to invest today to have your investment justc worth of cable service over 12 months
Answer #1

Let the amount invested be X

Monthly Cable bill P = $122

Interest rate on investment = 0.04/12 monthly

So, we need to calculate the PV of all the 12 payments that will be made in next 12 months

Hence, X = P/(1+r) + P/(1+r)2 +.... P/(1+r)12 = P [1 - (1+r)-n]/r = 122 (1 - (1+0.04/12)-12)/(0.04/12) = $1432.77

Hence, $1432.77 must be invested today at 4% to pay monthly cable bills of $122 for next 12 months